Suppose after charging a 3.1 µF capacitor to 14 V, you connect it to another already charged capacitor of 6.2 µF and 5 V. What voltage will your voltmeter read across those two capacitors connected in parallel?
Answer #1

Q1 = C1V1 = 3.1 x 14 = 43.4 uC

Q2 = C2V2 = 6.2 x 5 = 31 uC

total charge = 31 + 43.4 = 74.4 uC

net capacitance = 6.2 + 3.1 = 9.3 uF

potential = 74.4/9.3 = 8 V